OpenStudy (anonymous):

Solve for c. c plus two-fifths equals negative one-tenth c equals negative one-fifth c equals negative one-half c equals one-tenth c equals three-fifths

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

c + 2/5 = -1/10 (subtract 2/5 from each side) c = -1/10 - 2/5 (common denominator is 10) c = -1/10 - 4/10 c = - 5/10 = -1/2